OpenStudy (kaleidoscopicsink):

A particle is moving along the x-axis so that its position at t ≥ 0 is given by s(t) = (t)In(5t). Find the acceleration of the particle when the velocity is first zero. Help. I think the answer is 5e^2

OpenStudy (anonymous):

Velocity = derivative of position function. Acceleration = derivative of velocity function = second derivative of position function. So the rephrase the question: When is \(s'(t)=0\)? What's the value of \(s''\) at this point?

OpenStudy (anonymous):

You have \[s(t)=t\ln5t\implies s'(t)=\ln5t+1\implies s''(t)=\frac{1}{t}\] Find \(t\) such that \(\ln5t+1=0\) and plug it into the second derivative - pretty straightforward. How are you getting \(5e^2\)? You're almost right.

OpenStudy (kaleidoscopicsink):

Kinda guessing. Because it looked right. SO would that be just 5e then?

OpenStudy (anonymous):

Yes.

OpenStudy (kaleidoscopicsink):

Can you explain that to me a little bit?

OpenStudy (anonymous):

Which part? The actual computation, or why this is the way to do it?

OpenStudy (kaleidoscopicsink):

The actual computation. If that is no trouble.

OpenStudy (anonymous):

With \(s(t)=t\ln5t\), compute the first derivative via product rule: \[\begin{align*}\frac{\mathrm{d}s}{\mathrm{d}t}&=\frac{\mathrm{d}}{\mathrm{d}t}[t\ln5t]\\[1ex]&=\frac{\mathrm{d}}{\mathrm{d}t}[t]\ln5t+t\frac{\mathrm{d}}{\mathrm{d}t}[\ln5t]\\[1ex]&=\ln5t+t\frac{5}{5t}\\[1ex] &=\ln5t+1\end{align*}\] Set equal to \(0\) and solve for \(t\): \[\ln5t+1=0\implies 5t=\frac{1}{e}\implies t=\frac{1}{5e}\] Compute the second derivative: \[\begin{align*}\frac{\mathrm{d}^2s}{\mathrm{d}t^2}&=\frac{\mathrm{d}}{\mathrm{d}t}[\ln5t+1]\\[1ex] &=\frac{5}{5t}\\[1ex] &=\frac{1}{t}\end{align*}\] And finally, plug in \(t=\dfrac{1}{5e}\), so the value of the second derivative is simply \(5e\).

OpenStudy (kaleidoscopicsink):

Thank you so much. Can I have your assistance for one more question similar to this?

OpenStudy (anonymous):

You're welcome! I'm actually gearing up for a 4 hour drive in the next few minutes, so I'll be indisposed. There are plenty of other users with a calc proficiency, and calc questions tend to have a pretty high rate of getting a timely response.

OpenStudy (kaleidoscopicsink):

Thank you for your help. Have fun with your drive. cx